# Parallelism Meets Adaptiveness: Scalable Documents Understanding in Multi-Agent LLM Systems

### v3

@@ Abstract
Large language model (LLM) agents have shown increasing promise for collaborative task completion. However, existing multi-agent frameworks often rely on static workflows, fixed roles, and limited inter-agent communication, reducing their effectiveness in open-ended, high-complexity domains. This paper proposes a coordination framework that enables adaptiveness through three core mechanisms: dynamic task routing, bidirectional feedback, and parallel agent evaluation. The framework allows agents to reallocate tasks based on confidence and workload, exchange structured critiques to iteratively improve outputs, and—crucially—compete on high-ambiguity subtasks with evaluator-driven selection of the most suitable result. We instantiate these principles in a modular architecture and demonstrate substantial improvements in factual coverage, coherence, and efficiency over static and partially adaptive baselines. Our findings highlight the benefits of incorporating both adaptiveness and structured competition in multi-agent LLM systems.
@@ 2 Related Work
Multi-agent systems built on large language models (LLMs) have rapidly emerged as a powerful paradigm for solving complex, multi-stage tasks. Early frameworks such as AutoGPT and BabyAGI introduced single-agent recursive task planners with memory and subtask management, though they were limited by brittle, linear execution flows and lacked true inter-agent coordination [13, 17]. Subsequent systems like CAMEL [18] and CrewAI [11] introduced role-based delegation and dialogue-based collaboration among agents. LangGraph formalized agent workflows using graph structures [12], but did not incorporate mechanisms for feedback loops or task reassignment.
Recent meta-agent and hierarchical systems, such as MetaGPT [6] and Voyager [16], demonstrate improved planning, code synthesis, and environmental interaction. However, they typically assume static roles or unidirectional flows of control, lacking adaptiveness in task routing or feedback integration. The HuggingGPT framework [10] explores task orchestration by using ChatGPT to coordinate models on Hugging Face, yet still relies on fixed assignments. Generative Agents [8] illustrate long-term memory and simulation of social behaviors, providing inspiration for agent autonomy but not competitive refinement.
A more recent trend incorporates reflective or competitive coordination. For instance, systems like ChatDev [9] and AgentVerse [3] explore inter-agent negotiation and emergent behaviors. Others such as GameGPT [2] and DesignGPT [5] use LLM agents in creative or design-oriented domains, but often lack deep feedback-driven revision loops. Reflective multi-agent systems [1] combine agent memory with iterative critique, approaching the kind of meta-level coordination our system emphasizes.
Our work is also informed by broader surveys of multi-agent LLM collaboration [14, 7], which identify key challenges including error propagation, brittle delegation, and a lack of robustness under dynamic task flows. In contrast, our framework explicitly introduces dynamic routing, feedback-based revision, shared long-term memory, role self-optimization, and competitive evaluation—yielding a system that is not only collaborative but also adaptively reflective.
In a complementary domain, Wu et al. [15] explore competitive task scheduling for warehouse robots using reinforcement learning, demonstrating that concurrent multi-agent evaluation can improve efficiency even in physical task settings. Cross-domain applications of multi-perspective analysis and feedback-driven reconstruction, such as in civil infrastructure using deep learning and multiview stereo methods [4], further support the value of evaluation-oriented designs in complex systems. We build on this insight in the LLM space by introducing competitive agent evaluation to select the best output from parallel attempts.
@@ 3.2 Motivating Use Case
Consider a collaborative LLM-based system tasked with producing a technical report on a scientific topic. Such a system may involve:
  • Research agents that collect and summarize relevant academic sources.
  • Drafting agents that write specific sections of the report.
  • Evaluation agents that assess intermediate outputs and request revisions or improvements.
  • Parallel agents (optional) that generate competing drafts or analyses when ambiguity is high.
Without adaptive coordination, this team may suffer from redundant literature searches, inconsistencies between sections, or the propagation of factual errors. For example, if two research agents unknowingly retrieve the same source, effort is wasted; if an early factual mistake goes unchecked, it can taint all subsequent sections. These pitfalls motivate a more flexible coordination model that supports real-time feedback, dynamic task reallocation, shared memory of progress, and parallel trials of competing solutions when appropriate.
@@ 5.1 Parallel Execution and Selection Mechanism
Example Scoring Function $\mathcal{E}$
A composite score could be defined as:
$\mathcal{E}(o)=\alpha\cdot\text{Coherence}(o)+\beta\cdot\text{Factuality}(o)+\gamma\cdot\text{Relevance}(o)$
where $\alpha,\beta,\gamma$ are user-defined weights (e.g., $\alpha=0.3$, $\beta=0.4$, $\gamma=0.3$). These may be set by task domain or learned from historical outcomes.
@@ 5.2 Interaction Flow
At runtime, the orchestrator decomposes the task and assigns subtasks. Depending on context:
  • If the task is straightforward, it routes it to a single role agent.
  • If uncertainty or ambiguity is high, it triggers parallel execution and scoring.
  • Outputs are logged to shared memory, and any agent can retrieve them.
  • If feedback is issued, the orchestrator routes it back to the relevant agent or reassigns the task.
@@ 5.3 Modularity and Extensibility
The architecture supports plug-and-play expansion. One can introduce:
  • New role types (e.g., Visualizer Agent, Critique Agent)
  • Alternative memory backends (e.g., document DB vs. vector DB)
  • Custom scoring models (e.g., fine-tuned LLMs as evaluators)
Agents are loosely coupled via the shared memory and feedback bus, which supports parallelism and failure recovery. This makes the system scalable and adaptable to diverse task environments.
@@ 6.3 Ablation Study
6.4 Takeaways
  • Adaptive mechanisms reduce error propagation and improve synthesis quality.
  • Human-style review from an evaluator agent improves factuality and organization.
  • Dynamic routing avoids overload and allows agents to operate closer to their strengths.

### v4

@@ Abstract
Large language model (LLM) agents have shown increasing promise for collaborative task completion. However, existing multi-agent frameworks often rely on static workflows, fixed roles, and limited inter-agent communication, reducing their effectiveness in open-ended, high-complexity domains. This paper present a multi-agent coordination framework that improves the accuracy of Large Language Models (LLMs) in complex financial document analysis. Unlike existing frameworks that rely on static routing or linear workflows, our approach introduces Parallel Agent Evaluation, a mechanism where multiple agents compete on high-ambiguity subtasks. A centralized evaluator scores these parallel outputs based on factuality and coherence to select the optimal result. We evaluate this architecture on SEC 10-K filings, demonstrating a 27% improvement in compliance accuracy and a 74% reduction in revision rates compared to standard static baselines. These results validate that structured competition and dynamic routing significantly reduce hallucinations in high-stakes document understanding.
@@ 2 Related Work
Multi-agent systems built on large language models (LLMs) have rapidly emerged as a powerful paradigm for solving complex, multi-stage tasks. Early frameworks such as AutoGPT and BabyAGI introduced single-agent recursive task planners with memory and subtask management, though they were limited by brittle, linear execution flows and lacked true inter-agent coordination [torres2023auto, babyagi2023github]. Subsequent systems like CAMEL [zhu2023camel] and CrewAI [crewai2024github] introduced role-based delegation and dialogue-based collaboration among agents. LangGraph formalized agent workflows using graph structures [langgraph2024], but did not incorporate mechanisms for feedback loops or task reassignment.
Recent meta-agent and hierarchical systems, such as MetaGPT [hong2023metagpt] and Voyager [xu2023voyager], demonstrate improved planning, code synthesis, and environmental interaction. However, they typically assume static roles or unidirectional flows of control, lacking adaptiveness in task routing or feedback integration. The HuggingGPT framework [shen2023hugginggpt] explores task orchestration by using ChatGPT to coordinate models on Hugging Face, yet still relies on fixed assignments. Generative Agents [park2023generative] illustrate long-term memory and simulation of social behaviors, providing inspiration for agent autonomy but not competitive refinement.
A more recent trend incorporates reflective or competitive coordination. For instance, systems like ChatDev [qian2024chatdev] and AgentVerse [chen2024agentverse] explore inter-agent negotiation and emergent behaviors. Others such as GameGPT [chen2023gamegpt] and DesignGPT [ding2023designgpt] use LLM agents in creative or design-oriented domains, but often lack deep feedback-driven revision loops. Reflective multi-agent systems [bo2024reflective] combine agent memory with iterative critique, approaching the kind of meta-level coordination our system emphasizes.
Our work is also informed by broader surveys of multi-agent LLM collaboration [wang2023survey, lamalfa2025miss], which identify key challenges including error propagation, brittle delegation, and a lack of robustness under dynamic task flows. In contrast, our framework explicitly introduces dynamic routing, feedback-based revision, shared long-term memory, role self-optimization, and competitive evaluation—yielding a system that is not only collaborative but also adaptively reflective.
In a complementary domain, Wu et al. [wu2025warehouse] explore competitive task scheduling for warehouse robots using reinforcement learning, demonstrating that concurrent multi-agent evaluation can improve efficiency even in physical task settings. Cross-domain applications of multi-perspective analysis and feedback-driven reconstruction, such as in civil infrastructure using deep learning and multiview stereo methods [dan2024evaluation], further support the value of evaluation-oriented designs in complex systems. We build on this insight in the LLM space by introducing competitive agent evaluation to select the best output from parallel attempts.
@@ 3.2 Motivating Use Case
Consider a collaborative LLM-based system tasked with producing a technical report on a scientific topic. To ensure comprehensiveness and accuracy, the system orchestrates four distinct agent roles. Research Agents are responsible for collecting and summarizing relevant academic sources. Drafting Agents utilize these summaries to write specific sections of the report. Evaluation Agents assess intermediate outputs, requesting revisions or improvements where necessary. Finally, Parallel Agents are optionally deployed to generate competing drafts or analyses when ambiguity is high. Without adaptive coordination, this team may suffer from redundant literature searches or error propagation; however, our framework synchronizes these roles through shared memory and dynamic routing.
@@ 5.1 Parallel Execution and Selection Mechanism
5.2 Scoring Function Definition
To strictly quantify output quality, we employ a hierarchical scoring function $\mathcal{E}(o)$ driven by a Critic Agent. The score is calculated as a weighted sum of three normalized sub-metrics:
$\mathcal{E}(o)=w_{f}\cdot S_{fact}+w_{c}\cdot S_{coh}+w_{r}\cdot S_{rel}$ (1)
Where:
  • $S_{fact}$ (Factuality) is the ratio of claims in output $o$ supported by the retrieved context $C$, defined as $\frac{|supported\_claims|}{|total\_claims|}$.
  • $S_{coh}$ (Coherence) is the logical consistency score (010-1) derived from a chain-of-thought critique of the argument flow.
  • $S_{rel}$ (Relevance) measures the semantic cosine similarity between the output embedding and the query embedding.
In our experiments, we set weights to $w_{f}=0.5,w_{c}=0.3,w_{r}=0.2$ to prioritize factual accuracy, reflecting the high precision requirements of financial compliance tasks.
@@ 5.3 Interaction Flow
At runtime, the orchestrator decomposes the task and assigns subtasks based on the immediate context.
Standard Routing. If the task is assessed as straightforward, the orchestrator routes it directly to a single specialized role agent for immediate execution.
Parallel Execution. If uncertainty or ambiguity exceeds a predefined threshold, the orchestrator triggers parallel execution, spawning multiple agents to attempt the task simultaneously.
Memory Logging. All generated outputs are logged to the shared long-term memory, ensuring any agent can retrieve prior context regardless of the execution path.
Feedback Loop. If downstream agents issue feedback, the orchestrator routes the request back to the relevant agent or reassigns the task to a new agent better suited to address the critique.
@@ 5.4 Modularity and Extensibility
The architecture supports plug-and-play expansion, allowing the system to adapt to diverse domains. Extensible Roles allow for the introduction of new specialized agents, such as a Visualizer Agent or a dedicated Critique Agent. Flexible Memory Backends enable the swapping of storage solutions, supporting both document databases and vector stores depending on retrieval needs. Customizable Scoring permits the integration of domain-specific evaluation models, such as fine-tuned LLMs, to replace generic scoring functions. Agents remain loosely coupled via the shared memory and feedback bus, ensuring that these expansions do not disrupt existing workflows.
@@ 6 Case Study: Adaptive Coordination for 10-K Analysis
In addition to a static sequential baseline, we compared our framework against LangGraph’s standard multi-agent supervisor pattern [langgraph2024] as a state-of-the-art baseline. While the standard supervisor pattern improves over single-agent performance, our parallel evaluation mechanism demonstrated a further 14% improvement in compliance accuracy over the LangGraph baseline, specifically in high-ambiguity scenarios where single-path routing often fails to capture nuance.
@@ 6.3 Ablation Study
6.4 Key Findings
Our experiments on financial document analysis highlight three primary implications for multi-agent system design.
Mitigation of Error Propagation. Adaptive mechanisms, specifically bidirectional feedback, effectively sever the ”cascade of errors” common in static chains. By allowing downstream agents to reject low-quality inputs, we observed a 73% reduction in redundancy penalties, ensuring that early-stage hallucinations do not contaminate final reports.
Resilience via Structured Competition. The Parallel Agent Evaluation mechanism demonstrates that redundancy is not merely inefficient but essential for high-ambiguity tasks. By generating diverse interpretations of complex disclosures and selecting the optimal output via a scoring function, the system achieves a level of compliance accuracy (0.94) that single-path workflows cannot replicate.
Optimization of Agent Specialization. Dynamic task routing moves beyond simple load balancing to enable true role specialization. The system autonomously offloads technical legal parsing to compliance-focused agents while reserving summarization agents for narrative tasks, resulting in a workflow that is both 14% faster and contextually more robust than static assignments.
@@ 7 Discussion
7.1 Limitations and Scope
While our framework improves robustness in financial document understanding, it has several limitations. First, parallel agent evaluation increases inference cost and can add latency variance, making it less suitable for low-latency or cost-constrained settings. Second, the evaluator can introduce systematic bias if the scoring criteria are misaligned with downstream objectives, particularly for table-heavy or numeric-heavy sections. Third, our experiments focus on SEC 10-K filings; results may not directly generalize to domains with weaker document structure or less explicit grounding. Finally, our ambiguity detection and routing policies are heuristic rather than learned; developing learned policies from feedback is an important direction for future work.